# Constructing equivalences between fusion categories of quantum groups and of vertex operator algebras via quantum gauge groups

### v4

@@ 2. The main result
For part (b), we follow a suitable Drinfeld twist method originally used by Drinfeld for Drinfeld-Kohno theorem [25], [27], [28], and in our case for different categories, defined in the statement of the main result. This is the main part of the result. The first main step is the construction of a class of finite dimensional weak Hopf $C^{*}$-algebras $A_{W}({\mathfrak{g}},q)$ for all Lie types and positive integer levels, extending the former examples of [12] in the type $A$ case. They have unitary coboundary structures associated to the fusion category of quantum groups at roots of unity, associated to a specific linear fiber functor $W:{\mathcal{C}}({\mathfrak{g}},q)\to{\rm Vec}$, that emerges from the work of Wenzl [99]. Complete definitions of unitary coboundary structures on weak quasi-Hopf algebras are given in [13], and are motivated by the work of several authors, among them Drinfeld, Wenzl, Mack and Schomerus.
In Sect. 9 of [13] we prove two main abstract uniqueness results for associativity morphisms and braided symmetries in a semisimple pre-tensor category with a generating object. These are Theorem 9.13 and its extension Theorem 9.19. We explain here briefly how Theorem 9.19 applies to imply part (d) of Theorem 2.1.
The proof of parts (b), (c), and (d) is based on the general theory of semisimple weak quasi-Hopf algebras, weak Hopf algebras, and the theory of Drinfeld twist treated in Sections 5–8, and 10–18 in [13]. This abstract theory splits in two parts. Theory on which part (b), (c) of Theorem 2.1 is based is developed in Sections 5–8, and 10–17. The applicative part of the proof of part (b), (c) is developed in Sections 20–39. The abstract aspects that lead to the proof of part (d) of Theorem 2.1 are the main subject of Section 9.
For parts (b), (c) we develop an abstract analogue of Drinfeld-Kohno theorem 30.4 in [13], and apply it to the setting of fusion categories of quantum groups at suitable roots of unity and of module categories of affine vertex operator algebras at positive integer levels. We construct weak Hopf algebras $A_{W}({\mathfrak{g}},q)$ associated to the quantum group fusion category for all the Lie types in Section 32 in [13], by constructing a weak tensor structure on a linear functor from this category to the category of vector spaces, that emerges in [99] for the special tensor products described in (b), (c). The weak tensor structure is constituted by fusion intertwiners with respect to the action of $U_{q}({\mathfrak{g}})$.
@@
By application of our analogue of Drinfeld-Kohno theorem we obtain the structure of a weak quasi-Hopf algebra on the Zhu algebra $A(V_{{\mathfrak{g}}_{k}})$ associated to a corresponding affine vertex operator algebra $V_{{\mathfrak{g}}_{k}}$. Thus a natural weak quasitensor structure defined by fusion intertwiners with respect to the action of ${\mathfrak{g}}$ compatible with the fusion rules of modules of the vertex operator algebra. This induces the structure of a ribbon braided tensor category on the module category of $V_{{\mathfrak{g}}_{k}}$, via transport from the the Zhu algebra $A(V_{{\mathfrak{g}}_{k}})$.
@@
On the module category of $V_{{\mathfrak{g}}_{k}}$ we also have the ribbon braided tensor structure constructed by Huang and Lepowsky, and we need to compare the two structures.
@@
We discuss this transfer in Sect. 16 in [13] without the need of a weak quasi-tensor structure on ${\mathcal{F}}$, in abstract terms. One only needs an inverse linear equivalence ${\mathcal{S}}:{\rm Rep}(A)\to{\mathcal{C}}$ of the Tannakian linear equivalence ${\mathcal{E}}:{\mathcal{C}}\to{\rm Rep}(A)$. In this situation we may define a braided tensor category structure in ${\rm Rep}(A)$ from the braided tensor category structure in ${\mathcal{C}}$ with the property that ${\rm Rep}(A)$ is equivalent to ${\mathcal{C}}$ as a braided tensor category. The tensor product, associativity morphisms, braiding morphisms of ${\rm Rep}(A)$ are explicitly defined via the equivalence ${\mathcal{E}}$ and its inverse ${\mathcal{S}}$.
@@
Direct verification of coincidence of the two braided tensor structures on all the irreducible objects is more complicated due to multiplicity of fusion and non-semisimplicity of $U_{q}({\mathfrak{g}})$. To this aim rather than pursuing a closer analogy with the original Drinfeld-Kohno theorem for all objects, we develop two uniqueness theorems, Theorem 9.13 and its generalization Theorem 9.19. This generalization is made necessary by the actual conclusions reached in our work in the applicative sections on the comparison between special cases of the two associativity morphisms and their corresponding braided symmetries. Thus the uniqueness theorem 9.19 is formulated with suitable assumptions, where again in the application the abstract notion of ${\mathcal{V}}$-pre-associator of CFT-type for the fundamental representation $V$ is used in a central way both on the side of quantum groups and that of vertex operator algebras to verify coincidence of these special braid and associativity morphisms, after application of our Drinfeld-Kohno theorem. But this brief description of Theorems 9.13 and Theorem 9.19 put together with coincidence of the two braided tensor structures on special collections of objects proved in part (b), (c), is still not enough to describe our method of proof to complete the proof of part (d) and with it the proof of Theorem 2.1.
We need more that we describe in the following sections. The first next step is a somewhat complicated cohomological reduction work done in Theorems 9.13 and 9.19, where we use cohomology of the pentagon equation and its interplay with the hexagonal equations defining the braided symmetry for a general semisimple braided tensor category endowed with a generating object.
@@
Similarities with the original Drinfeld-Kohno theorem regard the treatment of special braiding morphisms, that is based as in Drinfeld proof on a twist given by a square root of the Drinfeld coboundary matrix. This approach is particularly useful for braiding on tensor products of the form $V_{\lambda}\otimes V$ or $V\otimes V_{\lambda}$, where multiplicity-free and structural fusion properties of $V$ allow for a direct verification of the equivalence between the two braided symmetries on these objects. Moreover in our case on special triples (or $n$-tuples) where two (or consecutive $n-1$-tuples) coordinates are fixed to $V$, verification of the equivalence of the two associativity morphisms similarly put in a nutshell reduces to the same properties of $V$. We do not use the KZ equation to this aim, or in any part of our work, other than the fact that the KZ equation was used by Drinfeld, Kazhdan and Lusztig, and by Huang and Lepowsky in the setting of affine vertex operator algebras, to construct the structure of a braided tensor category.
@@
Differently from the original Drinfeld-Kohno theorem, we do not study the problem of equivalence associativity morphisms in our version version of Drinfeld-Kohno theorem. We do this in Sect. 9 in [13] with the mentioned different ideas that appeal to duality properties of the fundamental representation $V$ with respect to the representation of the braid group in the tensor powers of $V$.
@@
The uniqueness Theorems 9.13 and 9.19 both apply under a central duality property of the fundamental representation $V$ with respect to the braid group representation in the quantum group tensor category, which is at the core of quantum group representation theory. The first instance of this property is Schur-Weyl duality for the general linear group. This duality says that the algebra of intertwining operators, from a tensor power of the vector representation $V$ of the general linear group to itself (also called centralizer algebras for the tensor powers of the vector representation) is generated by permutation operators [104].
@@
We need the property that the centralizer algebras $(V^{n},V^{n})$ with respect to the action of the Drinfeld-Jimbo quantum group $U_{q}({\mathfrak{g}})$ are generated by the natural representation of the braid group induced by the $R$-matrix. In our case $q$ is a root of unity as usually assumed corresponding to the unitary cases of the associated fusion category, specified in [99], [92], see also Def. 20.1, 21.2 in [13]. The tensor powers $V^{n}$ appearing in our setting are canonical choices of truncations of the full tensor powers of the quantum group $U_{q}({\mathfrak{g}})$ obtained by induction over $n$.
The construction of the iterated truncations $V^{n}$ of $U_{q}({\mathfrak{g}})$ at the root of unity is explained in [99] and is the starting point to construct our weak Hopf algebras in Sect. 32. The duality property between the truncated tensor powers $V^{n}$ of the fundamental representation and the representation of the braid group is needed as a further property to complete the proof of FKL theorem on coincidence between associativity morphisms and braiding morphisms on all pairs and triples and is used in our cohomological theorems 9.13 and 9.19. These theorems complete with cohomogical arguments what was proved up to that point in our work, coincidence of these structure maps on pairs and triples where one or two coordinated are fixed to the fundamental representation $V$. To do so, the way how our weak Hopf algebras are described in the application to the quantum group fusion category is also used.
@@
Returning to the question of whether the braid group representation generates the centralizer algebras $(V^{n},V^{n})$ for the truncated tensor powers of $U_{q}({\mathfrak{g}})$ at the root of unity, this question can be studied in the same way as for the case of a corresponding summand with $q$ not a root of unity, as in the truncated summand the root of unity property is not visible. This is explained among other things in Prop. 2.4 in [99], see also Prop 31.2 in [13] for the same result with a slightly more detailed proof, where an important role is played by the continuous curve described in that proposition, that passes through generic values of $q$. See also Section 3.5 in [99] for the special tensor products $V_{\lambda}\otimes V$ of representations of $U_{q}({\mathfrak{g}})$. Thus it suffices to have the duality property between the full algebras $(V^{\otimes n},V^{\otimes n})$ of $U_{q}({\mathfrak{g}})$ when $q$ is not a root of unity.
@@
In the application of Theorems 9.13 and 9.19, we work on the side of the quantum group fusion category and by the results above, our assumptions on braid group duality for the fundamental representation of our weak Hopf algebras $A_{W}({\mathfrak{g}},q)$ constructed in Sect. 32 in [13] are verified for the defining structure arising from Tannakian construction on the quantum group fusion category ${\mathcal{C}}({\mathfrak{g}},q))$ and Wenzl functor.
@@
Thus the fact that braid group duality for the fundamental representation $V$ was the relevant missing observation helpful to complete the proof of Theorem 2.1 in the previous version of [13], which led us to Theorems 9.13 and 9.19 as a completion of Lemma 7.10 in version 7 of [13], was for us a further important recognition of the relevance of the role of quantum groups. We refer the reader to the beautiful historical paper by V. F. R. Jones [64] for a general first introduction to the origin of quantum groups as solutions to the Yang-Baxter equation.
@@
We also would like to point out that our results of the categorical characterization of the class of unitary coboundary weak Hopf algebras that we obtain in Sect. 26, 27, 29 in [13] extends the notion of symmetric functors that plays a key role in Doplicher-Roberts theory [23], [24]. In Sect. 28 we explain the relation with the more familiar class of hermitian ribbon structures.
@@ 13. Outlook
 [^1] The author department digital devices were subject to a severe hacking attack for several months in 2021 which was recognized by authorities. The hacking severely damaged her intellectual property and health, and the hacked material has not been restituted. Related results have appeared in the papers [8], [9], [10]. These papers have been sued by the author to the Italian Procura della Repubblica between fall 2021 and spring 2023. Their claims, arguments, and reasoning essentially depend on [13], in-depth discussions between the author and S. Carpi that took place from 2018 to 2022 on the contents of an entire project including the topics of the above papers before his unilateral break of the collaboration in early 2023 and formal give up of his contribution, that consisted in a draft of a single section in 2018 based on [12], and yet unpublished joint work previously done by Sergio Ciamprone and the author (the first 60-page-long version of [13]), the original Finkelberg theorem, and work by Hans Wenzl in [99]. The author shared proofs of these attacks to several colleagues from 2022 to 2024.

### v5

@@ 4. The general plan of our proof
For part (b), we follow a suitable Drinfeld twist method originally used by Drinfeld for Drinfeld-Kohno theorem [22], [24], [25], and in our case for the stated categories. This step splits in various parts. The first main step is the construction of the class of finite dimensional unitary coboundary weak Hopf $C^{*}$-algebras $A_{W}({\mathfrak{g}},q)$ recalled in Theorem 2.4 for all the Lie types and positive integer levels. This construction extends with different methods the previos examples of [9] in the type $A$ case. The unitary coboundary structure of $A_{W}({\mathfrak{g}},q)$ naturally arises from the unitary structure of the fusion category of quantum groups at roots of unity, associated to the specific linear fiber functor $W:{\mathcal{C}}({\mathfrak{g}},q)\to{\rm Vec}$, that emerges from the work of Wenzl [99]. The second main step to prove (b) is the construction of the Drinfeld twist $F$ defined as a suitable square root of $\overline{R}$. A third step consists in the development of an abstract analogue of Drinfeld-Kohno theorem 29.4 in [10], that we apply to our twist composed with a quantization isomorphism described in [99] for the unitary structure of irreducible respresentations of the fusion category, between $A_{W}({\mathfrak{g}},q)$ and a corresponding quotient of $U({\mathfrak{g}})$ that identifies with the Zhu algebra $A(V_{{\mathfrak{g}}_{k}})$.
@@
By application of our analogue of Drinfeld-Kohno theorem we obtain the structure of a unitary coboundary weak quasi-Hopf algebra on the Zhu algebra $A(V_{{\mathfrak{g}}_{k}})$ associated to a corresponding affine vertex operator algebra $V_{{\mathfrak{g}}_{k}}$. Thus a natural weak quasitensor structure defined by fusion intertwiners with respect to the action of $\mathfrak{g}$} compatible with the fusion rules of modules of the vertex operator algebra. This induces the structure of a ribbon braided tensor category on the module category of $V_{{\mathfrak{g}}_{k}}$, via transport from the the Zhu algebra $A(V_{{\mathfrak{g}}_{k}})$. In particular, we get an independent construction of the associativity morphisms and proof of rigidity of ${\rm Rep}(V_{{\mathfrak{g}}_{k}})$, directly arising from the quantum group fusion category via the equivalence induced by the twist.
@@
Part (c) initiates the comparison between Huang-Lepowsky braided tensor structure on ${\rm Rep}_{\rm HL}(V_{{\mathfrak{g}}_{k}})$ and our structure on ${\rm Rep}_{\rm QG}(A(V_{{\mathfrak{g}}_{k}}))$ and the axioms of braided tensor equivalence is done in [10] by a direct comparison for the special braiding morphisms and the special associativity morphisms listed in (c) on the Zhu algebra, passing through (a). Up to this point we do not need any knowledge on the space of intertwiners in the two categories, the linear equivalence is an isomorphism between any two corresponding morphism spaces by the Drinfeld twist construction.
The necessary general theory of semisimple weak quasi-Hopf algebras, weak Hopf algebras, and the theory of Drinfeld twist treated in Sections 4–7, and 9–17 in [10]. This abstract theory splits in two parts. Theory on which part (b), (c) of Theorem 3.1 is based is developed in Sections 4–7, and 9–16. The applicative part of the proof of part (b), (c) is developed in Sections 19–38. The abstract aspects that lead to the proof of part (d) of Theorem 3.1 are the main subject of Section 9.
In Sect. 8 of [10] we explain two main abstract uniqueness results for associativity morphisms and braided symmetries in a semisimple pre-tensor category with a generating object. These are Theorem 8.13 and its extension Theorem 8.19 in [10]. In the subsequent sections we explain how Theorem 8.19 applies to establish a first step to part (d) of Theorem 3.1. These are the main new abstract results to obtain the complete braided tensor equivalence, and rely on key aspects of the discovery of quantum groups. In Sect. 11 we apply these theorems to complete to proof of part (d) of Theorem 3.1. In this case we work on the comparison between the two structures transported to the side of the quantum groups, $A_{W}({\mathfrak{g}},q)$, and we need knowledge on the generating property by the braid group between tensor powers of the fundamental representation, to apply our uniqueness theorem.
@@
We discuss this transfer in Sect. 15 in [10] without the need of a weak quasi-tensor structure on $\mathcal{F}$}, in abstract terms. One only needs an inverse linear equivalence ${\mathcal{S}}:{\rm Rep}(A)\to{\mathcal{C}}$ of the Tannakian linear equivalence ${\mathcal{E}}:{\mathcal{C}}\to{\rm Rep}(A)$. In this situation we may define a braided tensor category structure in ${\rm Rep}(A)$ from the braided tensor category structure in $\mathcal{C}$} with the property that ${\rm Rep}(A)$ is equivalent to $\mathcal{C}$} as a braided tensor category. The tensor product, associativity morphisms, braiding morphisms of ${\rm Rep}(A)$ are explicitly defined via the equivalence $\mathcal{E}$} and its inverse $\mathcal{S}$}.
@@
Direct verification of coincidence of the two braided tensor structures on all the irreducible objects is more complicated due to multiplicity of fusion and non-semisimplicity of $U_{q}({\mathfrak{g}})$. To this aim rather than pursuing a closer analogy with the original Drinfeld-Kohno theorem for all objects, we develop two uniqueness theorems, Theorem 8.13 and its generalization Theorem 8.19 in [10]. This generalization is made necessary by the actual conclusions reached in our work in the applicative sections on the direct comparison between special cases of the two associativity morphisms and their corresponding braided symmetries. Thus the uniqueness theorem 8.19 is formulated with suitable assumptions, where in the application the abstract notion of $\mathcal{V}$}-pre-associator of CFT-type for the fundamental representation $V$ is used in a central way both on the side of quantum groups and that of vertex operator algebras to verify coincidence of these special braid and associativity morphisms, after application of our Drinfeld-Kohno theorem. But this brief description of Theorems 8.13 and Theorem 8.19 put together with coincidence of the two braided tensor structures on special collections of objects proved in part (b), (c), is still not enough to describe our method of proof to complete the proof of part (d) and with it the proof of Theorem 3.1.
We need more that we describe in the following sections. The first next step is a somewhat complicated cohomological reduction work done in Theorems 8.13 and 8.19, where we use cohomology of the pentagon equation and its interplay with the hexagonal equations defining the braided symmetry for a general semisimple braided tensor category endowed with a generating object.
@@
Similarities with the original Drinfeld-Kohno theorem regard the treatment of special braiding morphisms, that is based as in Drinfeld proof on a twist given by a square root of the Drinfeld coboundary matrix. This approach is particularly useful for braiding on tensor products of the form $V_{\lambda}\otimes V$ or $V\otimes V_{\lambda}$, where multiplicity-free and structural fusion properties of $V$ allow for a direct verification of the equivalence between the two braided symmetries on these objects. Moreover in our case on special triples (or $n$-tuples) where two (or consecutive n1n-1-tuples) coordinates are fixed to $V$, verification of the equivalence of the two associativity morphisms similarly put in a nutshell reduces to the same properties of $V$. We do not use the KZ equation to this aim, or in any part of our work, other than the fact that the KZ equation was used by Drinfeld, Kazhdan and Lusztig, and by Huang and Lepowsky in the setting of affine vertex operator algebras, to construct the structure of a braided tensor category.
@@
Differently from the original Drinfeld-Kohno theorem, we do not study the problem of equivalence associativity morphisms in our version version of Drinfeld-Kohno theorem. We do this in Sect. 8 in [10] with the mentioned different ideas that appeal to duality properties of the fundamental representation $V$ with respect to the representation of the braid group in the tensor powers of $V$.
@@
The uniqueness Theorems 8.13 and 8.19 both apply under a central duality property of the fundamental representation $V$ with respect to the braid group representation in the quantum group tensor category, which is at the core of quantum group representation theory. The first instance of this property is Schur-Weyl duality for the general linear group. This duality says that the algebra of intertwining operators, from a tensor power of the vector representation $V$ of the general linear group to itself (also called centralizer algebras for the tensor powers of the vector representation) is generated by permutation operators [104].
@@
We need the property that the centralizer algebras $(V^{n},V^{n})$ with respect to the action of the Drinfeld-Jimbo quantum group $U_{q}({\mathfrak{g}})$ are generated by the natural representation of the braid group induced by the $R$-matrix. In our case $q$ is a root of unity as usually assumed corresponding to the unitary cases of the associated fusion category, specified in [99], [92], see also Def. 19.1, 20.2 in [10]. The tensor powers $V^{n}$ appearing in our setting are canonical choices of truncations of the full tensor powers of the quantum group $U_{q}({\mathfrak{g}})$ obtained by induction over $n$.
The construction of the iterated truncations $V^{n}$ of $U_{q}({\mathfrak{g}})$ at the root of unity is explained in [99] and is the starting point to construct our weak Hopf algebras in Sect. 32. The duality property between the truncated tensor powers $V^{n}$ of the fundamental representation and the representation of the braid group is needed as a further property to complete the proof of FKL theorem on coincidence between associativity morphisms and braiding morphisms on all pairs and triples and is used in our cohomological theorems 8.13 and 8.19. These theorems complete with cohomogical arguments what was proved up to that point in our work, coincidence of these structure maps on pairs and triples where one or two coordinated are fixed to the fundamental representation $V$. To do so, the way how our weak Hopf algebras are described in the application to the quantum group fusion category is also used.
@@
Returning to the question of whether the braid group representation generates the centralizer algebras $(V^{n},V^{n})$ for the truncated tensor powers of $U_{q}({\mathfrak{g}})$ at the root of unity, this question can be studied in the same way as for the case of a corresponding summand with $q$ not a root of unity, as in the truncated summand the root of unity property is not visible. This is explained among other things in Prop. 2.4 in [99], see also Prop 30.2 in [10] for the same result with a slightly more detailed proof, where an important role is played by the continuous curve described in that proposition, that passes through generic values of $q$. See also Section 3.5 in [99] for the special tensor products $V_{\lambda}\otimes V$ of representations of $U_{q}({\mathfrak{g}})$. Thus it suffices to have the duality property between the full algebras $(V^{\otimes n},V^{\otimes n})$ of $U_{q}({\mathfrak{g}})$ when $q$ is not a root of unity.
@@
In the application of Theorems 8.13 and 8.19, we work on the side of the quantum group fusion category and by the results above, our assumptions on braid group duality for the fundamental representation of our weak Hopf algebras $A_{W}({\mathfrak{g}},q)$ constructed in Sect. 31 in [10] are verified for the defining structure arising from Tannakian construction on the quantum group fusion category ${\mathcal{C}}({\mathfrak{g}},q))$ and Wenzl functor.
@@
Braid group duality for the fundamental representation $V$ was the relevant observation helpful to prove part (d) of Theorem 3.1 with respect to the previous parts of the same theorem, which could reach the equivalence only partially as stated in the previous parts (a)–(c). This property was indeed used to apply Theorems 8.13 and 8.19 in [10] and reach the full braided tensor equivalence for the stated Lie types. We refer the reader to the historical paper by V. F. R. Jones [61] for a general first introduction to the origin of quantum groups as solutions to the Yang-Baxter equation.
@@
We also would like to point out that our results of the categorical characterization of the class of unitary coboundary weak Hopf algebras that we obtain in Sect. 25, 26, 28 in [10] extends the notion of symmetric functors that plays a key role in Doplicher-Roberts theory [20], [21]. In Sect. 28 we explain the relation with the more familiar class of hermitian ribbon structures.
@@ 13. Outlook
Appendix A Historical remarks on rigidity in affine Lie algebra and vertex operator algebra categories
The original work by Finkelberg refers to the cited work in physics by Moore and Seiberg and in algebraic geometry by Beilinson, Feigin, and Mazur [4] for the rigid braided tensor structure on $\tilde{\mathcal{O}}_{\ell}$. As I learned from discussions with Huang, rigidity was not proved in [4], and there was no proof of rigidity using this approach.
By a private communication with Huang on the history of this problem, I learned that it finally turned out that the construction by Finkelberg described in [30] connecting $\tilde{\mathcal{O}}_{\ell}$ with Kazhdan and Lusztig structure inherited by $\tilde{\mathcal{O}}_{-\ell}$ could not suffice to derive rigidity of $\tilde{\mathcal{O}}_{\ell}$ from that of $\tilde{\mathcal{O}}_{-\ell}$. This implied that the proof of the equivalence $\tilde{\mathcal{O}}_{\ell}\to\tilde{\mathcal{O}}_{-\ell}$ was far from complete.
As later reported by Finkelberg in [31], it turned out that the completion of the proof of that proposition leading to rigidity of $\tilde{\mathcal{O}}_{\ell}$ and the equivalence $\tilde{\mathcal{O}}_{\ell}\to\tilde{\mathcal{O}}_{-\ell}$, involved the Verlinde formula as well (which became another main result in Finkelberg’s paper). This reflects also on Finkelberg composed equivalence $\tilde{\mathcal{O}}_{\ell}\to\tilde{\mathcal{O}}_{-\ell}\to{\mathcal{C}}({\mathfrak{g}},q)$, with Kazhdan-Lusztig equivalence that moreover still passes through the negative integer levels. This implies in particular that the theorem does not include a few exceptional cases, including ${\mathfrak{g}}=E_{8}$ and $k=2$, which correspond to cases not treated in Kazhdan-Lusztig theory.
Following the introductions in [51], [58], Finkelberg work [30], [31] may be reinterpreted as giving the construction of a braided tensor category structure on $\tilde{\mathcal{O}}_{\ell}$. One may also conclude that these papers give a second proof of rigidity and modularity of this category.
On the other hand, Finkelberg equivalence theorem is stated as Theorem 7.0.2 in [1] without proof. The braided tensor equivalence between $\tilde{\mathcal{O}}_{\ell}\to{\mathcal{C}}({\mathfrak{g}},q)$ obtained from the combined works by Finkelberg, Kazhdan and Lusztig is considered indirect by several authors.
Acknowledgements. The methodology and results presented here are the outcome of a long-term project initiated in 2016. The author wishes to clarify that S. Carpi joined the collaboration in January 2018 and participated in the joint development of a draft section based on the ongoing project, an early 60-page version of [10] v1. He formally withdrew his signature on December 19, 2022 with a letter to C.P. stating wave of intellectual property rights. The authors of the papers resulting from this project are the solely main contributors and responsible for the constructions and the entirety of the proofs presented. The methods in the project build upon the author’s previous joint work with S. Ciamprone (2016–2018) and M. V. Giannone (2020–2022).